# Diversifying by Intent in Recommender Systems

### v1

@@ Abstract.
It has become increasingly clear that recommender systems overly focusing on short-term engagement can inadvertently hurt long-term user experience. However, it is challenging to optimize long-term user experience directly as the desired signal is sparse, noisy and manifests over a long horizon. In this work, we show the benefits of incorporating higher-level user understanding, specifically user intents that can persist across multiple interactions or recommendation sessions, for whole-page recommendation toward optimizing long-term user experience. User intent has primarily been investigated within the context of search, but remains largely under-explored for recommender systems. To bridge this gap, we develop a probabilistic intent-based whole-page diversification framework in the final stage of a recommender system. Starting with a prior belief of user intents, the proposed diversification framework sequentially selects items at each position based on these beliefs, and subsequently updates posterior beliefs about the intents. It ensures that different user intents are represented in a page towards optimizing long-term user experience.
We experiment with the intent diversification framework on one of the world’s largest content recommendation platforms, serving billions of users daily. Our framework incorporates the user’s exploration intent, capturing their propensity to explore new interests and content. Live experiments show that the proposed framework leads to an increase in user retention and overall user enjoyment, validating its effectiveness in facilitating long-term planning. In particular, it enables users to consistently discover and engage with diverse contents that align with their underlying intents over time, thereby leading to an improved long-term user experience.
Recommender Systems, User Intents, Diversification, Long-Term User Experience
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Personalization
@@ 1. Introduction
However, it has become increasingly clear that focusing on optimizing short-term engagement can inadvertently lead to undesirable outcomes and hurt long-term user experience. For example, recommender systems that optimize click-through rates (CTR) may end up recommending clickbait contents that are of bad quality (Wang et al., 2021). They can also lead to pigeon-holing effects (Chaney et al., 2018; Nguyen et al., 2014) that prevent users from exploring new interests, resulting in a bad long-term user experience. Recognizing the drawbacks of short-term focused recommender systems, research on recommender systems has been shifting towards improving beyond short-term engagement in order to optimize long-term user experience on the platform.
Optimizing long-term user experience through recommender systems is challenging though, as the desired long-term outcome is usually sparse, noisy and manifests over a long time horizon (Wang et al., 2022). To tackle this challenge, advanced sequence modeling techniques such as Transformers (Kang and McAuley, 2018; Sun et al., 2019) and reinforcement learning (Chen et al., 2019) have been leveraged in recommender systems to account for a longer consumption history and plan several steps ahead. These approaches mostly still rely on modeling individual user-item interactions in the same fashion as the short-term focused recommender systems, and are black-box in nature. However, focusing on individual user-item modeling limits the recommender system’s planning horizon, as it is very hard to establish a causal relationship between a single recommendation item and the long-term user experience. In this work, we argue that a higher level understanding of the user’s intermediate needs and preferences is needed in order to provide the users with consistent recommendation experience and enable long-horizon planning.
The key observation that motivates our proposed framework is that user behaviors on the online platforms are largely driven by their underlying intents (Fishbein and Ajzen, 1977). These intents capture the latent causes or purposes driving specific decisions or actions from the user, such as making a purchase or consuming a particular item on the platform. These intents can vary not only across different users but also within the same user over time. For example, on a video recommendation platform, users with a learning intent may look for videos that help them acquire knowledge, information, or skills; those with an entertainment intent may be interested in videos that are funny, relaxing and recreational in nature. For the same user, their preferences may shift across different time and contexts: They might exhibit a learning intent on weekday mornings and an entertainment intent during weekend evenings or getaway trips. In addition, the most relevant content under the same intent can still vary across users or for the same user across different time and contexts. For example, a user with a learning intent might be interested in learning yoga in the morning and learning cooking in the afternoon. We therefore argue a crucial goal for recommendation platforms is to understand and cater to users’ real-time intents by recommending contents that align with their intents in a personalized and contextualized way, in order to improve their experience on the platform in the long term.
@@
As the first attempt to build towards intent based recommendation frameworks, we propose to ensure the system can accurately represent the diverse and evolving user intents through the assortment of items recommended on a given page in real time [^1] There are other alternative treatments to incorporate intents, which we will detail in a longer version that is out of scope of this paper.. Incorporating user intents in personalization has been primarily investigated within the context of search (Hu et al., 2011; Cheng et al., 2010), but remains largely under-explored for recommender systems. While an explicit user query can be mapped to a user intent in search; it is often missing in recommendation context, making it harder to identify the user intents, and letting alone to match recommendation items accordingly. To this end, we develop an intent prediction module based on user and context information, and then an intent diversification framework that is applied to the final stage of a recommender system, by adapting a greedy algorithm originally proposed for search results diversification (Agrawal et al., 2009). The diversification framework starts with a prior belief of user intents, represented by a probability distribution over different intents of the user, predicted by the intent prediction module. It sequentially selects items from the top position to the bottom, maximizing the probability of satisfying the user intent distribution at each step. After selecting each item at each position, it adjusts the posterior beliefs about user intents by assuming that the user is not interested in the previously placed items, thereby ensuring that a diverse set of intents are accurately represented in the final recommendations.
@@
We experiment with our proposed intent diversification framework on one of the world’s largest content recommendation platforms, serving billions of users everyday. In particular, we instantiate the framework by incorporating the user’s exploration intent, capturing their propensity to explore new interests and content. Live experiments on the platform show that our intent diversification framework enables users to consistently discover and engage with diverse contents that align with their underlying intents over time. In particular, when the users exhibit a higher propensity of exploration intent, our framework recommends more novel content that aligns with their preferences compared to users with lower propensity of exploration intent. Our framework leads to an increase in user retention, i.e., Daily Active Users (DAU), validating its effectiveness in optimizing long-term user experience.
@@ 2. Related Work
2.1. Optimizing Long-Term User Experience on Recommender Systems
Research on optimizing long-term user experience with recommender systems can be grouped into three main categories. The first category focuses on leveraging advanced sequence modeling techniques that allow the recommender systems to account for longer user history in modeling. Examples include Recurrent Neural Nets (RNNs) (Chen et al., 2019) and Transformers (Sun et al., 2019; Kang and McAuley, 2018) for recommender systems. The second category explores reinforcement learning (RL) approaches (Afsar et al., 2022; Kokkodis and Ipeirotis, 2021; Liu, 2022) that plan several steps ahead. The third category focuses on designing surrogate objectives that are predictive of the long-term outcome but easier to optimize (Athey et al., 2019; Yang et al., 2023; Simester et al., 2020; Wang et al., 2022).
These approaches still rely on modeling individual user-item interactions in the same fashion as the short-term focused recommender systems, by predicting the next item that will be consumed based on a sequence of past items that the user has interacted with. However, it is an intrinsically challenging problem to establish a causal relationship between a single recommendation item and the long-term user experience, as the desired long-term outcome is usually sparse, noisy and naturally manifests over a long horizon (Wang et al., 2022). As a result, existing techniques are inherently constrained by their item-level nature, which limits their planning horizon. In this work, we show that it is beneficial to go beyond item-level modeling and incorporate a higher-level understanding of the users in the ML models, in order to achieve long-horizon planning.
@@ 2.2. User Intent Modeling and Application
2.3. Diversification and Whole-Page Optimization
Industrial recommender systems started as two-stage framework (Covington et al., 2016): Candidate generation and ranking. The candidate generation stage provides broad personalization and narrows down millions of candidates to a few hundred. The ranking stage further selects a few dozen of items from the nominated candidates. Both stages rely on pointwise scores to select items that are of top relevance to the user. In other words, similarity among the selected items are ignored. Seeing a page of very similar items however can be a suboptimal recommendation experience for the users as it prevents them from exploring diverse interests (Chaney et al., 2018). Recognizing this challenge, state-of-art recommender systems in recent years adopt a three-stage framework by adding a diversification or whole-page optimization stage as the final stage of the recommender system (Wilhelm et al., 2018; Gao et al., 2023), to ensure that a diverse set of contents is represented as the final recommendations. This stage is a combinatorial optimization problem in nature as it needs to decide the optimal ordering of a set of candidates selected from the ranking stage. Several methods have been proposed over the years to reduce the complexity of the problem and achieve scalable diversification in large-scale systems, dating back to Carbonell’s work (Carbonell and Goldstein, 1998) which proposed to greedily add items to the recommendation list with maximal marginal relevance (MMR). Recent works on diversification and whole-page optimization include leveraging determinantal point processes (DPP) (Wilhelm et al., 2018), listwise modeling with recurrent or attentive sequence modeling (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019; Huang et al., 2020), and reinforcement learning (Zou et al., 2019). To our knowledge, current research on diversification and whole-page optimization still exclusively rely on modeling user-item interactions without considering user intents, indicating a missed opportunity. Our proposed intent diversification framework addresses this gap by integrating user intents into the final stage of a recommender system, to ensure that different user intents are accurately represented in the final recommendations.
@@ 3. Method
We first introduce the intent diversification framework, with a brief description of the intent prediction module at the end of the section.
@@ 3.1. Notations and Definitions
To introduce users’ dynamic intents into diversification, we denote the space of intents as $\mathcal{V}$, with $v$ representing a specific intent within this space. Let $\Pr(v|i)$ denote the prior belief of user $i$’s propensity to have intent $v$ for the current page, which could be derived from the output of a classification model trained on past user behaviors. Let $Q(j|i)$ represent the value function of item $j$ for user $i$, which is readily available in existing intent-agnostic recommendation systems. We adopt a probabilistic definition of $Q(j|i)$, where it denotes the probability that user $i$ enjoys item $j$. This probabilistic definition is chosen so that its complement, $1-Q(j|i)$, represents the probability that user $i$ does not enjoy item $j$. As we will see shortly, this formulation facilitates the development of a probabilistic framework for diversification, and is also widely adopted approach within search contexts (Agrawal et al., 2009).
@@ 3.2. Intent Diversification Algorithm
With the components defined and derived in Section 3.1, we are now ready to derive the intent diversification algorithm. Our method is motivated by the search diversification algorithm proposed by (Agrawal et al., 2009) for diversifying search results for ambiguous queries. In the context of search, a query could belong to multiple categories and can have more than one interpretation (Radlinski and Dumais, 2006). The authors propose to greedily diversify the search results by selecting one result at a time that maximizes the probability of satisfying an average user, assuming the previously presented results do not satisfy them. We bring such logic to the recommendation space by mapping a query to a user, and mapping the query’s category to the user’s intent. Specifically, by law of total expectation, the probability that item $j$ satisfies user $i$, considering all the possible intents on a page is equal to:
@@
In a search diversification framework such as (Agrawal et al., 2009), the item that maximizes Eq.(3) is selected as the top search result. However, it is not straightforward to directly apply such search diversification algorithm to the recommendation context. Unlike the search context, where relevance typically serves as the single objective, recommender systems involve multiple objectives summarized into a single quality score, $s_{ij}$, derived from various machine learning model outputs. As a result, this score often lacks a probabilistic interpretation. To integrate such non-probabilistic information into a probabilistic framework, we propose multiplying the quality score with the probabilistic intent-based prediction in Eq.(3). Specifically, at the first step of the intent diversification algorithm, we define the item at the top position, $j_{1}$ as
@@
Starting from the second position, the choice of which item to place at each subsequent position is influenced by the items already placed above that position. Imagine a user who examines the top 2 positions in order in making a decision whether to consume either item or quit. The objective of the recommender system is to maximize the value from either of the top 2 items. Intuitively, this implies that deciding what to put at the second position involves a counterfactual consideration, assuming that the user is not interested in the first item. Therefore, the second position should pick an item that is less likely to cover the intent that the first item has.
@@
Formally, the intent diversification algorithm described above is detailed in Algorithm 1 below. On a high level, the algorithm initiates with $R_{0}=\emptyset$, iteratively selecting one item $j_{m}$ at each step $m$ to be added to $R_{m-1}$ to formulate $R_{m}$ for $m=1,\ldots,M$, and concludes after $M$ steps at $R_{M}=(j_{1},\ldots,j_{M})$ which serves as the final intent-diversified recommendation results.
@@
Figure 1 shows an intuitive illustration of the proposed intent diversification algorithm, in particular the posterior intent update in Eq.(5) or equivalently, Line 11 in Algorithm 1. As depicted in the figure, user exploration intent starts with the highest propensity $p(Exploration)$ among the four intents (exploration, learning, gaming, fun). As a result, the top item selected is likely to be an exploratory item [^3] Note that it is not necessarily true that the top selected item is guaranteed to be an item from the intent with highest propensity. Instead, it also depends on the quality of the item according to Eq.(7).. Subsequently, the posterior belief on exploration intent drops because in the second position, we assume that the user is not interested in the first item, leading to a decreased posterior belief on the exploration intent. As a result, the relative importance of the rest of the intents (learning, gaming, fun) will increase, therefore more likely to show up in the subsequent items. Intuitively, by assuming that the user is not interested in previously placed items, this greedy and sequential diversification procedure ensures that all intents are well represented in the recommendation. By incorporating the multiplication of the quality score $s_{ij}$, it further ensures that the item selected at each position maintains high relevance to the users.
@@
3.3. User Exploration Intent
We now instantiate the framework to incorporate user exploration intent. In the recommender system literature, user exploration is defined as identifying unknown user interests or introducing users to new interests, where the latter is also known as serendipity (Herlocker et al., 2004). Chen et al. (2021) showed that user exploration and serendipity is tightly connected to long-term user experience on content recommendation platforms. Motivated by the previous findings in the literature, we choose user exploration intent as the intent for diversifying recommendation results toward optimizing long-term user experience on the platform.
In particular, we build a real-time machine learning model that predicts the propensity of the user having an exploration intent in the current page. It is built as a supervised deep learning model with the users’ past exploration behavior as labels. Specifically, the label for each page $t$ is defined as
(8) $y=\begin{cases}1,&\text{if the user consumes an exploration item on the page }t,\\ 0,&\text{otherwise},\end{cases}$
where an exploration item is defined as coming from a content provider with whom the user has not previously engaged. It is a page-level prediction, with the features including user behavior signals such as past consumption patterns, session-level features such as session length and average completion ratio in the current session, and contextual features such as time of the day and day of the week. We adopt a supervised binary classification model with cross-entropy loss. Figure 2 shows the architecture of the exploration intent model.
Figure 2. User exploration intent model.

### v3

# Beyond Item Dissimilarities: Diversifying by Intent in Recommender Systems

@@ Abstract.
It has become increasingly clear that recommender systems that overly focus on short-term engagement prevents users from exploring diverse interests, ultimately hurting long-term user experience. To tackle this challenge, numerous diversification algorithms have been proposed as the final stage of recommender systems. These algorithms typically rely on measures of item similarity, aiming to maximize the dissimilarity across items in the final set of recommendations. However, in this work, we demonstrate the benefits of going beyond item-level similarities by utilizing higher-level user understanding—specifically, user intents that persist across multiple interactions or recommendation sessions—in diversification. Our approach is motivated by the observation that user behaviors on online platforms are largely driven by their underlying intents. Therefore, recommendations should ensure that diverse user intents are accurately represented. While intent has primarily been studied in the context of search, it is less clear how to incorporate real-time dynamic intent predictions into recommender systems.
To address this gap, we develop a probabilistic intent-based whole-page diversification framework for the final stage of a recommender system. Starting with a prior belief of user intents, the proposed framework sequentially selects items for each position based on these beliefs and subsequently updates posterior beliefs about the intents. This approach ensures that different user intents are represented on a page, towards optimizing long-term user experience.
We experiment with the intent diversification framework on YouTube, the world’s largest video recommendation platform, serving billions of users daily. Live experiments on a diverse set of intents show that the proposed framework increases Daily Active Users (DAU) and overall user enjoyment, validating its effectiveness in facilitating long-term planning. Specifically, it enables users to consistently discover and engage with diverse content that aligns with their underlying intents over time, leading to an improved long-term user experience.
Recommender Systems, User Intent, Diversification, Long-Term User Experience
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3690624.3709429ccs: Information systems Recommender systemsccs: Information systems Personalization
@@ 1. Introduction
However, it has become increasingly clear that focusing on optimizing short-term engagement can inadvertently lead to undesirable outcomes and hurt long-term user experience. In particular, this can lead to pigeon-holing and echo-chamber effects (Chaney et al., 2018; Nguyen et al., 2014) that prevent users from exploring new interests, resulting in a bad long-term user experience. To tackle this challenge, numerous exploration and diversification algorithms have been proposed to improve the novelty and diversity of the recommendation results (Carbonell and Goldstein, 1998; Chen et al., 2021; Wilhelm et al., 2018). These algorithms rely item-level notions such as novelty (Herlocker et al., 2004; Chen et al., 2021) or pairwise similarities (Carbonell and Goldstein, 1998; Ziegler et al., 2005), aiming to maximize the dissimilarity across items in the final set of recommendations.
In this work, we demonstrate the benefits of going beyond item-level similarities by utilizing a higher-order user understanding—specifically, user intents that persist across multiple interactions or recommendation sessions—in diversifying recommendation results. This is motivated by the observation that user behaviors on the online platforms are largely driven by their underlying intents (Fishbein and Ajzen, 1977). These intents can vary not only across different users but also within the same user over time. Therefore, an ideal diversification algorithm should account for the different intents that a user may have. In particular, it should diversify not only across topics but also across the different intents of the user, in order to provide coherent recommendation results and improve user experience in the long term.
@@
Incorporating user intents in diversification has been primarily investigated within the context of search (Hu et al., 2011; Cheng et al., 2010), but remains largely under-explored for content recommendation platforms. To bridge this gap, in this work we develop an intent diversification framework that is applied to the final stage of a recommender system, by adapting a greedy algorithm originally proposed for search results diversification (Agrawal et al., 2009). The diversification framework starts with a prior belief of user intents, represented by a probability distribution over different intents of the user, predicted by the intent prediction module. It sequentially selects items from the top position to the bottom, maximizing the probability of satisfying the user intent distribution at each step. After selecting each item at each position, it adjusts the posterior beliefs about user intents by assuming that the user is not interested in the previously placed items, thereby ensuring that a diverse set of intents are accurately represented in the final recommendations.
@@
We experiment with our proposed intent diversification framework on YouTube, the world’s largest video recommendation platform, serving billions of users everyday. We instantiate the framework with a diverse set of user intents. Live A/B experiments on YouTube demonstrate that our intent diversification framework enables users to consistently discover and engage with diverse content that aligns with their underlying intents over time. Our framework leads to significant improvements in business metrics, including user retention (i.e., Daily Active Users, DAU) and overall user enjoyment, validating its effectiveness in optimizing long-term user experience.
@@ 2. Related Work
2.1. Diversification in Recommender Systems
Industrial recommender systems started as two-stage framework (Covington et al., 2016): Candidate generation and ranking. The candidate generation stage provides broad personalization and narrows down millions of candidates to a few hundred. The ranking stage further selects a few dozen of items from the nominated candidates. Both stages rely on pointwise scores to select items that are of top relevance to the user. In other words, similarity among the selected items are ignored. However, seeing a page of very similar items however can be a suboptimal recommendation experience for the users as it may lead to echo-chamber and pigeon-holing effects, preventing the users from exploring diverse interests (Chaney et al., 2018; Ge et al., 2020). Recognizing this challenge, state-of-art recommender systems in recent years adopt a three-stage framework by adding a diversification or whole-page optimization stage as the final stage of the recommender system (Wilhelm et al., 2018; Gao et al., 2023), to ensure that a diverse set of contents is represented as the final recommendations (Ziegler et al., 2005). This stage is a combinatorial optimization problem in nature as it needs to decide the optimal ordering of a set of candidates selected from the ranking stage. Several methods have been proposed over the years to reduce the complexity of the problem and achieve scalable diversification in large-scale systems, dating back to Carbonell’s work (Carbonell and Goldstein, 1998) which proposed to greedily add items to the recommendation list with maximal marginal relevance (MMR). Recent works on diversification and whole-page optimization include leveraging determinantal point processes (DPP) (Wilhelm et al., 2018), listwise modeling with recurrent or attentive sequence modeling (Zhuang et al., 2018; Ai et al., 2018; Pei et al., 2019; Huang et al., 2020), graph theories(Adomavicius and Kwon, 2011), and reinforcement learning (Zou et al., 2019).
To our knowledge, most diversification algorithms to date still exclusively rely on item-level similarity and dissimilarity measures, lacking a higher-level understanding of user intents. This indicates a missed opportunity for providing a coherent recommendation experience based on user intents that supports long-term user engagement and satisfaction. Our proposed intent diversification framework addresses this gap by integrating user intents into the final stage of a recommender system, ensuring that different user intents are accurately represented in the final recommendations towards a coherent and improved long-term user experience.
@@ 3.1. Notations and Definitions
To introduce users’ dynamic intents into diversification, we denote the space of intents as $\mathcal{V}$, with $v$ representing a specific intent within this space. Let $\Pr(v|i)$ denote the prior belief of user $i$’s propensity to have intent $v$ for the current page, which is the output of a classification model trained on past user behaviors as described below in Section 3.2. Let $Q(j|i)$ represent the value function of item $j$ for user $i$, which is readily available in existing intent-agnostic recommendation systems. We adopt a probabilistic definition of $Q(j|i)$, where it denotes the probability that user $i$ enjoys item $j$. This probabilistic definition is chosen so that its complement, $1-Q(j|i)$, represents the probability that user $i$ does not enjoy item $j$. As we will see shortly, this formulation facilitates the development of a probabilistic framework for diversification, and is also widely adopted approach within search contexts (Agrawal et al., 2009).
@@
3.2. User Intent Modeling
An important input to our proposed intent diversification algorithm is the real-time intent probabilities $\Pr(v|i)$. We propose a supervised ML model $f^{v}(\cdot;\theta)$, parameterized by $\theta$, to generate a personalized, contextualized, and real-time probability that user $i$ has intent $v$ in the current recommendation page. The input features, denoted by $x$, include user behavior signals such as past consumption patterns, session-level features such as session length and average completion ratio in the current session, and contextual features such as time of the day and day of the week. [^2] For business compliance reasons, we omit the full feature list here.
Identifying a user’s intent without directly asking her is a significant challenge. To address this, we define the intent label by leveraging behavior signals on the page. Specifically, we let the label $y^{v}=1$ if the consumer consumes an item $j$ that aligns intent $v$ in the current recommendation page, i.e. $v\in\mathcal{V}_{j}$, and 0 otherwise. For example, we can define the user on a particular recommendation page to have an exploration intent if she consumes an exploration item on the page, where an exploration item is defined as coming from a content creator with whom the user has not previously engaged. Training the ML model entails minimizing the following cross-entropy loss (Hastie et al., 2009) on the training data $\mathcal{D}$:
(3) ^$\hat{\mathcal{L}}(\theta)=-\sum_{(x,y^{v})\in\mathcal{D}}\sum_{v\in\mathcal{V}}y^{v}\log f(x;\theta).$
In terms of the model architecture for $f^{v}(\cdot;\theta)$, we opt for a multi-layer neural network with a Deep & Cross Network (DCN) (Wang et al., 2017) component for efficiently capturing the interactions among different features. Figure 1 shows the architecture of the intent model.
Figure 1. User Intent Model.
@@ 3.3. Intent Diversification Algorithm
With the components defined and derived in Section 3.1 and 3.2, we are now ready to derive the intent diversification algorithm. Our method is motivated by the search diversification algorithm proposed by (Agrawal et al., 2009) for diversifying search results for ambiguous queries. In the context of search, a query could belong to multiple categories and can have more than one interpretation (Radlinski and Dumais, 2006). The authors propose to greedily diversify the search results by selecting one result at a time that maximizes the probability of satisfying an average user, assuming the previously presented results do not satisfy them. We bring such logic to the recommendation space by mapping a query to a user, and mapping the query’s category to the user’s intent. Specifically, by law of total expectation, the probability that item $j$ satisfies user $i$, considering all the possible intents on a page is equal to:
@@
In a search diversification framework such as (Agrawal et al., 2009), the item that maximizes Eq.(4) is selected as the top search result. However, it is not straightforward to directly apply such search diversification algorithm to the recommendation context. Unlike the search context, where relevance typically serves as the single objective, recommender systems involve multiple objectives summarized into a single quality score, $s_{ij}$, derived from various machine learning model outputs. As a result, this score often lacks a probabilistic interpretation. To integrate such non-probabilistic information into a probabilistic framework, we propose to multiply the quality score with the probabilistic intent-based prediction in Eq.(4). Specifically, at the first step of the intent diversification algorithm, we define the item at the top position, $j_{1}$ as
@@
Starting from the second position, the choice of which item to place at each subsequent position is influenced by the items already placed above that position. Imagine a user who examines the top 2 positions in order to make a decision whether to consume either item or quit. The objective of the recommender system is to maximize the value from either of the top 2 items. Intuitively, this implies that deciding what to put at the second position involves a counterfactual consideration, assuming that the user is not interested in the first item. Therefore, the second position should pick an item that is less likely to cover the intent that the first item has.
@@
Formally, the intent diversification algorithm described above is detailed in Algorithm 1 below. On a high level, the algorithm initiates with $R_{0}=\emptyset$, and iteratively selects one item $j_{m}$ at each step $m$ to be added to $R_{m-1}$ to formulate $R_{m}$ for $m=1,\ldots,M$. It concludes after $M$ steps at $R_{M}=(j_{1},\ldots,j_{M})$, which serves as the final intent-diversified recommendation results.
@@
Figure 2 shows an intuitive illustration of the proposed intent diversification algorithm. As an example, user exploration intent starts with the highest propensity $p(Exploration)$ among the four intents (exploration, learning, gaming, fun). As a result, the top item selected is likely to be an exploratory item. [^3] Note that it is not necessarily true that the top selected item is guaranteed to be an item from the intent with highest propensity. Instead, it also depends on the quality of the item according to Eq.(8). Subsequently, the posterior belief on exploration intent drops because in the second position, we assume that the user is not interested in the first item, leading to a decreased posterior belief on the exploration intent. As a result, the relative importance of the rest of the intents (learning, gaming, fun) will increase, therefore more likely to show up in the subsequent items. Intuitively, by assuming that the user is not interested in previously placed items, this greedy and sequential diversification procedure ensures that all intents are well represented in the recommendation. The multiplication of the quality score $s_{ij}$ further ensures that the item selected at each position maintains high relevance to the users.